OpenStudy (anonymous):

Particular solution to ODE y'' -2y' +5y = e^2t * (1-tsin(3t))

OpenStudy (anonymous):

I already solved the part after distributing. e^2t - e^2t *t*sin(3t) so now I need help solving -e^2t*t*sin(3t)

OpenStudy (kainui):

Can you show me what you've done so far in trying to solve it?

OpenStudy (anonymous):

So far, I used superposition to split up the particular term into e^2t and -e^2t*t*sin(3t). I have already solved the homogeneous case (setting right hand side to =0) and solving for y''-2y'+5y=e^2t, I get Xp1 = e^2t * 1/5 I have never encountered 3 terms and that is why I am slightly stumbled.

OpenStudy (kainui):

Oh ok, well what methods do you know about, I want to help you do it with a method you're comfortable with since I know of several ways you can attack this thing.

OpenStudy (kainui):

Undetermined coefficients I'm assuming is what you're doing?

OpenStudy (anonymous):

Yes that would be the most ideal method. I am just having trouble with this since its 3 terms and I never had an example of this before.

OpenStudy (alekos):

Hi Silen. I can help you with this, but it would take some time. How quickly do you need a solution?

OpenStudy (anonymous):

4 hours would be fine. If you could just tell me how to get started on the initial undetermined coefficient that would be a great start.

OpenStudy (anonymous):

You have a homogeneous part and a nonhomogeneous part to this ODE. Have you arrived at the solution to the homogeneous part?

OpenStudy (anonymous):

Oh my mistake, you posted the solution to the nonhomogeneous part where you only consider \(e^{2t}\)... Sorry!

OpenStudy (anonymous):

For the remaining part, since \(e^{2t}t\sin 3t\) isn't linearly dependent relative to either the homogeneous solution pair (\(y_h=e^t(\cos 2t+\sin2t)\)) as well as the first particular solution (\(y_p=\dfrac{1}{5}e^{2t}\)), you can set the remaining guess solution to something like \(y_p=e^{2t}(At+B)\sin3t\).

OpenStudy (anonymous):

Actually, a slight adjustment: Since the degree of this ODE is 2 and a first order derivative is included, you may need to consider a cosine term as well.

OpenStudy (anonymous):

@sithsandgiggles Ok so basically e^2t*(At+B)sin(3t)+ e^2t(Ct+D)cos(3t) ?

OpenStudy (anonymous):

@sithsandgiggles Ok so basically e^2t*(At+B)sin(3t)+ e^2t(Ct+D)cos(3t) ?

OpenStudy (anonymous):

That's what I would try, yep

OpenStudy (anonymous):

Taking those derivatives might be a nightmare but hopefully it works.

OpenStudy (alekos):

Siths is right on the mark. I'm going through the 1st and 2nd derivative of that large expression but it's proving very tedious. e^2t*(At+B)sin(3t)+ e^2t(Ct+D)cos(3t) still working on it

OpenStudy (alekos):

I've had no access to the site for quite some time due to some server issue.

OpenStudy (alekos):

It was a nightmare!